The numerator of a certain fraction is 5 more than the denominator. If 34 is added to the fraction, the result is 3. Find the fraction.

Answer by richwmiller(17219)

x=y+5,
x/y+34=3
can't be
you copied something wrong.
I bet it should be 3/4 not 34
x=y+5,
x/y+3/4=3
x = 9, y = 4